# Socratic Reasoning Improves Positive Text Rewriting

### v1

@@ 4.3 Evaluating Cognitive Reframers
Transfer strength defines how well negative sentiment is turned positive. To measure it, we first use a finetuned RoBERTa model to evaluate the sentiment scores of the original and reframed thoughts. Then, we use the pairwise difference of sentiment scores between the original and reframed thoughts and report the average, denoted by (Pos). We report the number of samples with an increase in positivity with respect to the original thought as accuracy. A decrease in the score from the original thought is considered a failed case (Acc).
Our aim is to reframe a thought but also preserve the original meaning and not drastically change it. We rely on the following automatic metrics which have been widely used in the text reframing literature: BLEU Papineni et al. (2002) and BLEURT Sellam et al. (2020). Following all three datasets we consider Ziems et al. (2022); Maddela et al. (2023); Sharma et al. (2023), we report the BLEU scores for each pair of original and reframed thoughts. Since it has been shown that BLEU scores do not always correlate well with human judgements on semantic similarity, we also use a BERT-based BLEU variation, BLEURT-20 which is trained on synthetic samples to get accurate semantic similarity scores. We report the average over a dataset. Each score is a value between 0 and 1, ranging from no to complete semantic similarity.
@@ 5.1 Main Results
Table 3 shows that using Socratic rationales consistently outperforms both vanilla finetuning and prompting strategies on all datasets. We show an example in Table 4 that highlights how the positive rewritings appear much more thoughtful, because the Socratic rationale forces the model to explicitly reason about the rewriting process. It is noteworthy that Socratic reasoning improves both the content preservation as well as transfer strength capabilities of smaller models beyond the performance of ChatGPT, despite the model being multiple times larger than LLaMa 7B and Mistral 7B.
Across all datasets, we observe that Mistral performs better in content preservation while LLaMA performs better in the sentiment of the reframed thought. This could be due to the different pretraining datasets used for training the models. We observe that using Socratic reasoning improves BLEU scores by almost 2 points on average when compared to finetuning without Socratic rationales. Distilling Socratic reasoning into smaller models helps to preserve the original meaning and improve the positive sentiment transfer. This is desirable for users and practitioners alike, because more positive sentiment will likely benefit users more and content preservation ensures that the reframing stays relevant for them.
@@ 5.2 Socratic Rationales are Informative
Similar to previous work for evaluating free-text rationales Chen et al. (2023a), we use the pretrained GPT-Neo 2.7B Gao et al. (2020) for computing REV. We report the average REV metric over each sample for all datasets. A REV value greater than zero implies that the Socratic rationale provides new, additional information for supporting the reframed thought. In contrast, a value less than zero implies the rationale does not support the reframe. As reported in Figure 4, we observe that all three datasets have REV values greater than zero, signifying the usefulness of enriching the cognitive reframing task with Socratic rationales.
Table 5 shows the generated rationales for each of the dataset we consider. In particular, POSREF has the highest informativeness for the reframed thoughts. We hypothesize two main reasons for this. First, POSREF does not contain any extra information (situations or personas) like the other two datasets where the information gets spread out through a large number of tokens. Second, samples in POSREF are sourced from Reddit and more representative of the training data of GPT-Neo than the crowdsourced data in the other two datasets.
Method |
Reframed Thought |
Mistral-FS |
I will improve. |
Mistral-CoT |
I will improve and learn to be better. |
Mistral-FT |
I will learn from the feedback and grow. |
Mistral-SoC |
It is normal to feel disappointed. I can use this experience to learn and grow. |
@@ 6.1 Evaluating Reframed Thoughts
While previous works in cognitive reframing Maddela et al. (2023); Sharma et al. (2023) rate the reframed thoughts on a Likert-scale with dimensions like fluency, helpfulness and readability, we instead consider that the main goal of such reframing systems is to help users overcome unwanted thoughts. Therefore, we conduct pairwise comparisons between 100 randomly sampled reframes generated by our proposed SocraticReframe finetuning strategy vs either ChatGPT or a baseline model (that does not use the Socratic rationales). Two computer science graduate students conducted the task. We report win rate as the number of times a reframe generated by the Socratic model is preferred by an annotator over the baseline reframe.
@@ 6.2 Evaluating Socratic Rationales
To evaluate the synthetically generated rationales, we consider two key attributes: whether they represent the use of Socratic questioning in a clinically meaningful way and the general helpfulness of the questioning to overcome negative thoughts.
Following Braun et al. (2015), we use the Socratic Questioning Scale (SQS) to evaluate the quality of generated rationales. SQS contains five questions rated on a Likert scale, which evaluate the degree to which the Socratic method was used in a given snippet. The final score is reported as the sum of the Likert ratings of the five questions. Since SQS was originally developed for assessing therapy session transcripts, we only include the three questions which are relevant for evaluating the use of the Socratic Method in a given text. The three questions evaluate the following dimensions for the rationale: Open-endedness, Context and Relevance to the situation. The full questions are reported in Appendix C.1. Additionally, we consider helpfulness as a metric to assess whether or not the generated rationale is generally helpful in overcoming negative thoughts. We ask the annotators to rate the generated rationales on a 3-point Likert scale.
Three computer science graduate students annotated 100 randomly sampled Socratic rationales across the three datasets. Figure 6 shows that the rationales were rated favorably across all criteria. We obtain an Intra Class Correlation (ICC) Shrout and Fleiss (1979) of 0.806 for helpfulness and 0.601 for the SQS scale. SQS appears to be particularly challenging to evaluate by the raters and judgements on the use of the Socratic Method may be subjective. Our observations align with works in psychology that report similar agreement scores for the SQS scale Braun et al. (2015). Additionally, we observe a Pearson correlation coefficient of 0.72 () between the helpfulness and SQS ratings, indicating that when the human raters evaluate a rationale as containing high-quality Socratic questioning, they also tend to deem it more helpful for dealing with the negative thought.
@@ 7 Conclusion & Future Work
In this work, we show that text rewriting models for cognitive reframing can be improved by using Socratic rationales to verbalize the reframing process. Because such rationales can be costly to obtain from human annotators, we use synthetically generated rationales to augment three open-source datasets for text rewriting. We then train models to verbalize the rationale before generating the reframing. We show that this leads to substantial improvements to finetuned open-sources LLMs like LLaMa-2 7B and Mistral 7B when compared to training without rationales and strong prompt-based baselines. We confirm our results with an extensive human evaluation that shows high ratings for the synthetically generated rationales and a clear preference for reframings generated with our method. By releasing our code and data we hope to enable future research, for example, on incorporating the different question types that are annotated, or using the data in therapist training.
@@
Acknowledgements
This work was supported by the DYNAMIC center, which is funded by the LOEWE program of the Hessian Ministry of Science and Arts (Grant Number LOEWE 1 16 519 03 09 001 0009 98). This research work has also been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
@@ Appendix A Experimental details
We few-shot prompt GPT-4 with the default parameters from the OpenAI Python package. 333https://github.com/openai/openai-python We use three few-shot exemplars. Following similar works on synthetic data generation with LLMs Macina et al. (2023a), we use temperature sampling with and no top-k truncation for generating the reframes. For the fine-tuning experiments, following Dettmers et al. (2024), we apply LoRA to all linear layers; training for 5 epochs with AdamW, a batch size of 8 and set the learning rate to 5e-4. All other hyperparameters take the default value in the configuration. We use NVIDIA A100 80GB for the training.
@@ C.1 Evaluating Socratic Rationales
We plan to release the generated rationales to complement the original datasets under the Apache license.444https://www.apache.org/licenses/LICENSE-2.0.
C.2 Evaluating Reframes
@@
Given the context and original negative thought, which reframed thought do you prefer (A vs B)?
@@ Appendix D Qualitative Examples
COGREF |
Situation: Having problems coming up with words to write essay. Negative Thought: I will get a bad grade. Reframed Thought: I’m feeling anxious about this assignment. I can use skills to calm myself, try my best on the essay, ask for any appropriate help, and whatever grade I get I will know that I tried my hardest. |
@@
Q: You’re worried about doing badly on this assignment. How have you done on past essays? A: I’ve gotten good grades, sometimes. But this one feels different. Q: Do you think getting a bad grade on this specific essay means you’re generally bad at writing or at this subject overall? A: Well, I don’t think so, but I’m unsure about this topic. Q: What evidence is leading you to believe that you will get a bad grade? A: I’m having trouble coming up with points to write about. Q: If a friend was in your position, what advice would you give them? A: I would probably say to not panic and just give it their best shot. Q: What would happen if you got a bad grade on this essay? How would that affect your overall academic achievement? A: One bad grade isn’t going to ruin my overall performance I guess. Q: Can using some support like talking to your teacher or a study group help you in getting through this? A: That might be a good idea. Q: So given these factors, do you still think it’s a definite that you will get a bad grade? A: No, I suppose it’s not definite. Q: So, what will you do about your essay now? A: I think I’ll start by focusing on writing a draft based on what I know and then ask for some help to see how I can improve it. |
PATREF |
Persona: my favorite band is rise against. my favorite team is the raptors. i usually go to bed late. i like sports. i am studying to be a dentist. Negative Thought: I am about to graduate and become a Dentist, but that means nothing. Anyone can do that. Reframed Thought: I am about to graduate and become a Dentist. Although it is possible for anyone to become a dentist, it is very difficult, and I am proud of myself. |
@@
Q: It sounds like you’re doing something remarkable, yet you downplay its value. Do you genuinely believe becoming a dentist means nothing? A: It doesn’t seem special when everyone else is doing similar things. Q: Do you think everyone could become a dentist? A: Maybe not everyone, but a lot of people could. Q: It required you to put in several years of hard study and practice, didn’t it? A: Yes, I did work hard. Q: How many people do you think are willing or able to do what you did? A: I’m not sure, probably not that many. Q: So wouldn’t that mean that your achievement is indeed quite significant and not something anyone could do? A: I guess so. Q: Isn’t it possible then that your achievement of becoming a dentist is both valuable and hard-earned? A: Yes, it’s starting to look that way. Q: How do you feel about your upcoming graduation now? A: I’m starting to recognize its importance and feel a bit proud. |
POSREF |
Negative Thought: This physics homework is the death of me every week. It’s so hard. And then I have another physics exam tomorrow night. Reframed Thought: Physics is challenging, but is worth it. I have an exam tomorrow, which I hope to do well on. |
@@
Q: It sounds quite challenging, is Physics always this difficult for you or only certain aspects? A: Certain parts, like quantum mechanics. Q: So, is it accurate to state that the entire subject is difficult, or just the quantum mechanics part? A: I suppose just the quantum mechanics part. Q: Can you give me an example of a physics problem or concept you’ve been able to understand quite well? A: Yeah, I’ve been doing well in classical mechanics. Q: Do you think that possessing skills in one area of physics could potentially aid in understanding other areas, including quantum mechanics? A: Maybe. But quantum mechanics seems very different. Q: If classical mechanics and quantum mechanics are two aspects of the same discipline, isn’t there a possibility that understanding one could help with the other? A: It’s a possibility, I guess. Q: Do you think saying that the physics homework is the death of you is a fair and accurate statement? A: Probably not, it’s just really frustrating sometimes. Q: What if you thought about the physics homework as a challenge rather than a death sentence? Could that change how you approach it? A: Maybe. I could try. Q: Now, how do you feel about the physics exam tomorrow? A: I’m nervous, but maybe if I look at it as a challenge, it might feel less overwhelming. |

### v2

@@ 4.3 Evaluating Cognitive Reframers
Transfer strength defines how well negative sentiment is turned positive. To measure it, we first use a finetuned RoBERTa model to evaluate the sentiment scores of the original and reframed thoughts. Then, we use the pairwise difference of sentiment scores between the original and reframed thoughts and report the average, denoted by (Pos). We report the number of samples with an increase in positivity with respect to the original thought as accuracy. A decrease in the score from the original thought is considered a failed case (Acc). Sharma et al. (2023) show that people tend to prefer more empathetic reframes over overly positive reframes. Hence, we also report Emp as the difference between empathy scores of original and reframed thoughts. The scores are computed using a pretrained RoBERTa empathy classifier. Our aim is to reframe a thought but also preserve the original meaning and not drastically change it. We rely on the following automatic metrics which have been widely used in the text reframing literature: BLEU Papineni et al. (2002) and BLEURT Sellam et al. (2020). Following all three datasets we consider Ziems et al. (2022); Maddela et al. (2023); Sharma et al. (2023), we report the BLEU scores for each pair of original and reframed thoughts. Since it has been shown that BLEU scores do not always correlate well with human judgements on semantic similarity, we also use a BERT-based BLEU variation, BLEURT-20 which is trained on synthetic samples to get accurate semantic similarity scores. We report the average over a dataset. Each score is a value between 0 and 1, ranging from no to complete semantic similarity.
@@ 5.1 Main Results
Table 3 shows that using Socratic rationales consistently outperforms both vanilla finetuning and prompting strategies on all datasets. We show an example in Table 4 that highlights how the positive rewritings appear much more thoughtful, because the Socratic rationale forces the model to explicitly reason about the rewriting process. We empirically check whether the generated reframes are close in length to the ground truth reframes for all datasets. We observe that the average length of generated reframes with our best performing model on all datasets is only 5 tokens longer than the ground truth, on average. It is noteworthy that Socratic reasoning improves both the content preservation as well as transfer strength capabilities of smaller models beyond the performance of ChatGPT, despite the model being multiple times larger than LLaMa 7B and Mistral 7B. Across all datasets, we observe that Mistral performs better in content preservation while LLaMA performs better in the sentiment of the reframed thought. This could be due to the different pretraining datasets used for training the models. We also observe that our method consistently generates more empathetic reframes which is generally preferred by people Sharma et al. (2023). We note that both positivity and empathy are lower for POSREF than other datasets, possibly because POSREF contains tweets which are known to be less positive in nature Sokolova et al. (2017). We observe that using Socratic reasoning improves BLEU scores by almost 2 points on average when compared to finetuning without Socratic rationales. Distilling Socratic reasoning (from GPT-4) into smaller models helps to preserve the original meaning and improve the positive sentiment transfer. This is desirable for users and practitioners alike, because more positive sentiment will likely benefit users more and content preservation ensures that the reframing stays relevant for them.
Method | Reframed Thought |
Mistral-FS | I will improve. |
Mistral-CoT | I will improve and learn to be better. |
Mistral-FT | I will learn from the feedback and grow. |
Mistral-SoC | It is normal to feel disappointed. I can use this experience to learn and grow. |
@@ 5.2 Socratic Rationales are Informative
Similar to previous work for evaluating free-text rationales Chen et al. (2023a), we use the pretrained GPT-Neo 2.7B Gao et al. (2020) for computing REV. We report the average REV metric over each sample for all datasets. A REV value > 0 suggests Socratic rationales support reframing by providing additional information, while < 0 indicates otherwise. In Figure 4, all datasets have REV > 0, highlighting Socratic rationale’s utility in enhancing reframing tasks. Table 5 shows the generated rationales for each of the dataset we consider. In particular, POSREF has the highest informativeness for the reframed thoughts. We attribute this to two main factors: First, POSREF lacks additional context like situations or personas, causing token dispersion where more information can lead to low values of REV. Second, POSREF samples from Reddit align better with GPT-Neo’s training data than the crowdsourced data in the other datasets.
5.3 Socratic Rationales progressively lead to a better reframe
We partition Socratic rationales into four quarters sequentially, each representing 25% of the data, and calculate average sentiment scores for each quarter. Figure 5 reveals a consistent trend across all datasets: intermediate answers lead to increasingly positive questions, suggesting that they aid in iterative reframing.
@@ 6.1 Evaluating Reframed Thoughts
Prior cognitive reframing studies Maddela et al. (2023); Sharma et al. (2023) use Likert scales for assessing aspects like fluency and readability. Yet, our focus lies on helping users overcome unwanted thoughts. Li et al. (2024) show that pairwise human preferences are well-defined. In a similar setup as Sharma et al. (2023), we compare 100 randomly selected SocraticReframe generated reframes against ChatGPT or a baseline model without Socratic rationales. After consenting to participate, two computer science graduate students were recruited to conduct the comparisons. The win rate indicates a preference for Socratic reframes over the baseline by the raters. The participants were asked to select the reframed thought they find most relatable, helpful and memorable - a criteria for good reframes defined in Sharma et al. (2023).
@@ 6.2 Evaluating Socratic Rationales
To evaluate the synthetically generated rationales, we consider two key attributes: whether they represent the use of Socratic questioning in a clinically meaningful way and the general helpfulness of the questioning to overcome negative thoughts. Following Braun et al. (2015), we use the Socratic Questioning Scale (SQS) to evaluate the quality of generated rationales. SQS contains five Likert-scale questions evaluating the use of the Socratic method in a snippet. The final score is the sum of these ratings. Originally developed for therapy session transcripts, we include only three relevant questions for text evaluation: Open-endedness, Context, and Relevance.333We note that in our evaluation criteria, “relevant” questions aim to tackle extrinsic hallucinations (generated text cannot be verified given the source) and “contextual” questions aim to tackle intrinsic hallucinations (generated text contradicts the source text), following the definitions from Ji et al. (2023). The full questions are reported in Appendix C.1. Additionally, we consider helpfulness as a metric to assess whether or not the generated rationale is generally helpful in overcoming negative thoughts. Annotators rate the generated rationales on a 3-point Likert scale.
Expert Evaluation
We recruit 14 mental health experts and clinical psychology graduate students for evaluating the quality of the Socratic rationales. We verify the quality of randomly sampled Socratic rationales with a clinical expert before conducting the study with psychology students. The annotation task was deemed exempt from an ethics review by an institutional review board. Amidei et al. (2019) suggest the use of Cronbach’s for checking the evaluation reliability in natural language generation tasks. We observe acceptably high values for helpfulness (0.67), diversity (0.69), reflectiveness (0.75) and relevance (0.65) of the questions in the rationale. We observe that raters validate the synthetically generated Socratic rationales as helpful to reframe the original thought. More than 50% of the raters also responded affirmatively on whether they are willing to use LLMs for cognitive reframing. More details are found in Section C.2.
Human Evaluations
Three computer science graduate students annotated 100 randomly sampled Socratic rationales from the datasets. Figure 7 shows favorable ratings across all criteria. We obtained an Intra Class Correlation (ICC) of 0.806 for helpfulness and 0.601 for the SQS scale, indicating that while the SQS is challenging to evaluate, helpfulness is generally agreed upon by non-experts. These findings align with similar agreement scores in psychology Braun et al. (2015). Additionally, a Pearson correlation coefficient of 0.72 () between helpfulness and SQS ratings suggests that better Socratic questioning correlates with greater helpfulness in addressing negative thoughts.
@@ 7 Conclusion & Future Work
In this work, we show that text rewriting models for cognitive reframing can be improved by using Socratic rationales to verbalize the reframing process. By releasing our code and data we hope to enable future research, for example, on incorporating the different question types that are annotated, or using the data in therapist training.
@@ Appendix A Experimental details
We few-shot prompt GPT-4 with the default parameters from the OpenAI Python package. 444https://github.com/openai/openai-python We use three few-shot exemplars. Following similar works on synthetic data generation with LLMs Macina et al. (2023a), we use temperature sampling with and no top-k truncation for generating the reframes. For the fine-tuning experiments, following Dettmers et al. (2024), we apply LoRA to all linear layers; training for 5 epochs with AdamW, a batch size of 8 and set the learning rate to 5e-4. All other hyperparameters take the default value in the configuration. We use NVIDIA A100 80GB for the training.
@@ C.1 Evaluating Socratic Rationales
We plan to release the generated rationales to complement the original datasets under the Apache license.555https://www.apache.org/licenses/LICENSE-2.0.
C.2 Validation with experts
We recruit psychology students for our study. The students were given credits in return of their annotation efforts. The average age of participants was 24.2 with 80% participants being females. The participants were shown three diverse samples from across the datasets and asked to rate the helpfulness and SQS, in the same setting as described in Section C.1. Additionally, participants were asked to answer the following question with either "Yes", "No" or "Not sure":
Would you use an LLM to help you restructure negative thoughts?
This question assesses the willingness and openness of psychology students towards incorporating Large Language Models (LLMs) into their therapeutic practices. Responses provide insights into whether LLMs are perceived as viable tools for cognitive restructuring, indicating their potential acceptance within psychological interventions.
We observe that only 7% of the participants did not want to use LLMs for restructuring negative thoughts with 55% responding affirmatively and the rest being unsure. We plan to conduct a more large-scale annotation study with more experts and non-experts in the future.
C.3 Evaluating Reframes
@@
Given the context and original negative thought, which reframed thought do you find more relatable, helpful and memorable (A vs B)?
@@ Appendix D Qualitative Examples
COGREF | Situation: Having problems coming up with words to write essay. Negative Thought: I will get a bad grade. Reframed Thought: I’m feeling anxious about this assignment. I can use skills to calm myself, try my best on the essay, ask for any appropriate help, and whatever grade I get I will know that I tried my hardest. | Q: You’re worried about doing badly on this assignment. How have you done on past essays? A: I’ve gotten good grades, sometimes. But this one feels different. Q: Do you think getting a bad grade on this specific essay means you’re generally bad at writing or at this subject overall? A: Well, I don’t think so, but I’m unsure about this topic. Q: What evidence is leading you to believe that you will get a bad grade? A: I’m having trouble coming up with points to write about. Q: If a friend was in your position, what advice would you give them? A: I would probably say to not panic and just give it their best shot. Q: What would happen if you got a bad grade on this essay? How would that affect your overall academic achievement? A: One bad grade isn’t going to ruin my overall performance I guess. Q: Can using some support like talking to your teacher or a study group help you in getting through this? A: That might be a good idea. Q: So given these factors, do you still think it’s a definite that you will get a bad grade? A: No, I suppose it’s not definite. Q: So, what will you do about your essay now? A: I think I’ll start by focusing on writing a draft based on what I know and then ask for some help to see how I can improve it. |
@@
PATREF | Persona: my favorite band is rise against. my favorite team is the raptors. i usually go to bed late. i like sports. i am studying to be a dentist. Negative Thought: I am about to graduate and become a Dentist, but that means nothing. Anyone can do that. Reframed Thought: I am about to graduate and become a Dentist. Although it is possible for anyone to become a dentist, it is very difficult, and I am proud of myself. | Q: It sounds like you’re doing something remarkable, yet you downplay its value. Do you genuinely believe becoming a dentist means nothing? A: It doesn’t seem special when everyone else is doing similar things. Q: Do you think everyone could become a dentist? A: Maybe not everyone, but a lot of people could. Q: It required you to put in several years of hard study and practice, didn’t it? A: Yes, I did work hard. Q: How many people do you think are willing or able to do what you did? A: I’m not sure, probably not that many. Q: So wouldn’t that mean that your achievement is indeed quite significant and not something anyone could do? A: I guess so. Q: Isn’t it possible then that your achievement of becoming a dentist is both valuable and hard-earned? A: Yes, it’s starting to look that way. Q: How do you feel about your upcoming graduation now? A: I’m starting to recognize its importance and feel a bit proud. |
@@
POSREF | Negative Thought: This physics homework is the death of me every week. It’s so hard. And then I have another physics exam tomorrow night. Reframed Thought: Physics is challenging, but is worth it. I have an exam tomorrow, which I hope to do well on. | Q: It sounds quite challenging, is Physics always this difficult for you or only certain aspects? A: Certain parts, like quantum mechanics. Q: So, is it accurate to state that the entire subject is difficult, or just the quantum mechanics part? A: I suppose just the quantum mechanics part. Q: Can you give me an example of a physics problem or concept you’ve been able to understand quite well? A: Yeah, I’ve been doing well in classical mechanics. Q: Do you think that possessing skills in one area of physics could potentially aid in understanding other areas, including quantum mechanics? A: Maybe. But quantum mechanics seems very different. Q: If classical mechanics and quantum mechanics are two aspects of the same discipline, isn’t there a possibility that understanding one could help with the other? A: It’s a possibility, I guess. Q: Do you think saying that the physics homework is the death of you is a fair and accurate statement? A: Probably not, it’s just really frustrating sometimes. Q: What if you thought about the physics homework as a challenge rather than a death sentence? Could that change how you approach it? A: Maybe. I could try. Q: Now, how do you feel about the physics exam tomorrow? A: I’m nervous, but maybe if I look at it as a challenge, it might feel less overwhelming. |